OpenStudy (anonymous):

Please help me solve this! I have done it over like 50 times!

OpenStudy (anonymous):

\[1+\sqrt{3x+3}=\sqrt{7x+2}\] I am super confused, i keep doing it by squaring both sides and i get x=.5 but it should be x=2

OpenStudy (perl):

Did you get this far? ( 1 + sqrt(3x+3))^2 = (sqrt(7x+2))^2 ( 1 + sqrt(3x + 3))*( 1 + sqrt(3x + 3)) = 7x + 2

OpenStudy (anonymous):

Oops, I think that is what I have been doing wrong. I thought i was just supposed to square them and it would look like this\[1+3x+3=7x+2\]

OpenStudy (perl):

Original equation: 1 + sqrt(3x+3) = sqrt(7x+2) square both sides ( 1 + sqrt(3x+3))^2 = (sqrt(7x+2))^2 ( 1 + sqrt(3x + 3))*( 1 + sqrt(3x + 3)) = 7x + 2 1*1 + sqrt(3x+3)+ sqrt(3x+3) + (sqrt(3x+3))^2 = 7x + 2 1 + 2*sqrt(3x+3) + 3x+3 = 7x + 2

OpenStudy (anonymous):

Hey @perl think you could tell me? I want to know just in case i get another problem like this.

OpenStudy (anonymous):

I wanted to know if 1 was the reason this happened, it must have gotten deleted.

OpenStudy (jhannybean):

\[1+\sqrt{3x+3}=\sqrt{7x+2}\]\[(1+\sqrt{3x+3})^2 = (\sqrt{7x+2})^2\]\[\sf \text{recall that} ~(a+b)^2 = a^2 +2ab+b^2\]\[1+2\sqrt{3x+3}+3x+3=7x+2\]\[2\sqrt{3x+3}+3x+4=7x+2\]\[2\sqrt{3x+3}=4x-2\]\[\sqrt{3x+3}=2x-1\]\[(\sqrt{3x+3})^2=(2x-1)^2\]\[3x+3=4x^2-2(2x)(-1)+1\]\[3x+3=4x^2+4x+1\]

OpenStudy (perl):

Original equation: 1 + sqrt(3x+3) = sqrt(7x+2) square both sides ( 1 + sqrt(3x+3))^2 = (sqrt(7x+2))^2 ( 1 + sqrt(3x + 3))*( 1 + sqrt(3x + 3)) = 7x + 2 1*1 + sqrt(3x+3)+ sqrt(3x+3) + (sqrt(3x+3))^2 = 7x + 2 1 + 2*sqrt(3x+3) + 3x+3 = 7x + 2 2*sqrt(3x+3) + 4 + 3x = 7x + 2 2*sqrt(3x+3) = 4x - 2 sqrt(3x+3) = 2x - 1 now square both sides (sqrt(3x+3))^2 = (2x-1)^2 3x+3 = (2x-1)^2 3x + 3 = (2x-1)*(2x-1) 3x+3 = 4x^2 - 2x - 2x + 1 3x+3 = 4x^2 - 4x + 1 4x^2 - 4x - 3x + 1 - 3 = 0 4x^2 - 7x - 2 = 0 (4x + 1) ( x - 2) = 0 x = -1/4 , x = 2 plug both solutions into original equation

OpenStudy (perl):

:)

OpenStudy (anonymous):

Thanks guys, this helps a lot.

OpenStudy (jhannybean):

Let me check with you too:\[3x+3=4x^2-4x+1\]\[4x^2-7x-2=0\]\[x^2-7x-8=0\]\[(x-8)(x+1)=0\]\[\left(x-\frac{8}{4}\right)\left(x+\frac{1}{4}\right)=0\]\[(x-2)(4x+1)=0\]\[x-2 \implies x=2\]\[4x+1 \implies x=-\frac{1}{4}\]\[\checkmark\]